# Quantum in Biology, Quantum for Biology, and Biology for Quantum: Mapping the Evidence and the Road Ahead

### 30 Apr 2026 v1

@@ III Quantum in Biology
FIG. 2: Overview of selected key topic areas in quantum in biology. The panels schematically represent DNA replication and proton transfer, enzyme catalysis and hydrogen tunneling, olfaction and receptor-level vibrational or electron-transfer hypotheses, DNA charge transport through base-stacked pathways, vision through retinal photoisomerization in rhodopsin, radical-pair magnetoreception, excitonic dynamics in photosynthesis, and proposed photophysical or collective effects in microtubules. The figure is illustrative and does not imply that all topics have the same evidential status. Figure generated using ChatGPT, with scientific content reviewed by the authors.
@@ III.3.4 Cellular respiration
Mitochondrial respiration is the chain of reactions in which electrons from metabolic substrates pass through redox cofactors in the respiratory complexes of the inner mitochondrial membrane. The free energy released by this stepwise electron transfer is coupled to proton translocation across the membrane, generating the electrochemical gradient used for ATP synthesis. The most plausible quantum contributions in this context are tunneling corrections to electron- and proton-transfer steps, since these processes naturally involve barriers and donor–acceptor distances for which quantum effects can become rate-relevant, particularly when coupled to protein reorganization and gating motions [156, 225, 88, 89, 90, 146, 27, 1]. More broadly, contemporary mitochondrial biology increasingly treats mitochondria as dynamic signaling organelles rather than passive ATP-producing units, reinforcing the importance of linking microscopic transport and redox models to cellular outputs and higher-level phenotypes [174, 226]. Evidence is strongest at the level of mechanistic modeling, computational identification of plausible transfer pathways, and biochemical constraints on distances and energetics, whereas proposals involving long-lived coherence or entanglement remain speculative and are only weakly constrained by direct in vivo measurements [85, 59, 163]. Decisive studies should therefore identify perturbations that alter tunneling likelihood while minimally affecting classical driving forces and reorganization, and then connect these microscopic changes to macroscopic outputs using orthogonal biochemical validation, with careful separation of changes in coupling from changes in the redox landscape and membrane potential.
III.3.5 Chiral induced spin selectivity in biological electron transfer
@@ III.6 Section synthesis
The quantum-in-biology landscape separates less by biological domain than by evidential maturity. The most secure cases are microscopic quantum-chemical contributions that are tightly constrained by kinetics and mechanism, especially enzymatic hydrogen tunneling and related nonadiabatic proton- or electron-transfer chemistry. An intermediate band contains areas such as radical-pair magnetoreception, photosynthetic exciton and vibronic dynamics, DNA charge transport, DNA repair photochemistry, and respiration-related transfer processes, where the mechanistic frameworks are physically serious but the case for biological necessity remains conditional and perturbation-dependent. A more speculative band contains system-level extrapolations, including strong microtubule-wide physiological claims, Posner-cluster neurobiology, and physical interpretations of quantum-information formalisms in genetic or neural information processing, where decisive tests are still missing or not yet sufficiently discriminating [129, 142, 97, 108, 71]. A compact topic-by-topic summary of this section is provided in Appendix Table 2.
@@ IV Quantum for Biology
FIG. 3: Overview of representative directions in quantum for biology. The panels illustrate quantum simulation of molecular and spin-dynamical systems, quantum-assisted molecular design and drug discovery, quantum machine learning for bioinformatics, nitrogen-vacancy sensing, quantum-enhanced NMR and magnetic-resonance methods, and quantum imaging or spectroscopy. The figure summarizes tool classes discussed in the section; it does not imply established quantum advantage for each application. Figure generated using ChatGPT, with scientific content reviewed by the authors.
@@ V Biology for Quantum
FIG. 4: Overview of representative directions in biology for quantum. The panels schematically show DNA origami scaffolds, virus-templated materials, DNA-templated wiring, biohybrid excitonic or polaritonic platforms, fluorescent protein qubits, biomolecular or cellular quantum-biosensing concepts, bioinspired transport, and quantum or quantum-inspired neural-network architectures. These examples illustrate ways biological structures or principles may support quantum technologies, with evidential strength depending on device-level benchmarks, reproducibility, and comparison to non-biological alternatives. Figure generated using ChatGPT, with scientific content reviewed by the authors.
@@ V.4.1 Noise-robust transport and interference inspired by photosynthesis
V.4.2 Brain-inspired quantum and quantum-inspired neural computation
@@ V.5 Section synthesis
Biology for quantum is strongest when the biological contribution is not merely illustrative but difficult to replace by conventional design strategies, and when that contribution is measured through explicit device, fabrication, sensing, or algorithmic benchmarks. DNA origami and related self-assembly platforms already provide concrete enabling capabilities for nanoscale positioning and fabrication. Biohybrid polaritonic systems and biomolecular spin platforms are scientifically credible emerging directions, but they still require stronger benchmarks for stability, reproducibility, and device-level performance. By contrast, broader brain-inspired or quantum-like architectures should presently be treated as hypothesis-generating and design-motivating unless they can be tied to explicit hardware, sensing, or algorithmic advantages under matched comparisons [121, 128, 68, 2, 3, 198]. A compact topic-by-topic summary of this section is provided in Appendix Table 4.
@@ VI Evolution and Future Perspectives
A useful discipline across this interface is to require every strong claim to answer five questions explicitly: What specific quantum resource is being invoked? What is the minimal mechanistic or algorithmic model that generates distinctive predictions? What is the strongest classical alternative under the same conditions? What perturbation, benchmark, or matched comparison would discriminate between them? And what biological or engineering output would count as success? Claims that cannot yet be mapped onto this framework are often best presented as hypotheses or design directions rather than as established results. This caution is particularly important in adjacent areas such as photobiomodulation and light-responsive mitochondrial signaling, where reported biological effects are clearly of interest, even though the full mechanistic chain remains incompletely resolved and is unlikely to be captured by a single explanatory model [192, 133, 12]. Similarly, work on ultra-weak photon emission highlights the importance of distinguishing experimentally established phenomena from stronger claims about coherence or nonclassicality, for which reliable evidence remains limited [51].
@@ VII Acknowledgments
During the preparation of this manuscript, GPT-assisted tools were used in a limited editorial role to support clarity, organization, and language refinement. In addition, the schematic images included in the manuscript were generated using GPT-assisted image tools. All scientific content, literature selection, interpretation, figure selection, and final wording were independently reviewed and approved by the authors. The authors take full responsibility for the accuracy, integrity, and originality of the manuscript.
@@ Appendix A Summary tables across the map
Direction Topic cluster Present status Most efficient upgrade path Quantum in Biology Molecular quantum chemistry and transformation Most secure where microscopic quantum-chemical effects are tightly constrained by kinetics and mechanism, especially enzymatic hydrogen tunneling and related PCET chemistry Link microscopic transfer mechanisms more tightly to full catalytic cycles, polymerase-resolved outcomes, and perturbation-resolved biological phenotypes Quantum in Biology Sensory transduction Strongest for photon detection and radical-pair spin chemistry; olfaction remains more theoretical and receptor-level evidence is limited Use receptor-, protein-, and organism-level perturbations tied directly to sensory outputs, with controls for classical binding, adaptation, and transduction effects Quantum in Biology Excited-state, redox, and spin transfer Strong physicochemical or ultrafast evidence in several systems, but functional necessity under biological conditions remains conditional and perturbation-dependent Use dephasing-, detuning-, redox-, spin-, or photochemical perturbations tied directly to yield, signaling, repair, or bioenergetic output Quantum in Biology Macromolecular transport and collective effects Experimentally and theoretically active, especially for microtubule photophysics and spin-sensitive tubulin effects, but not yet established as physiological quantum mechanisms Use independent replication, standardized preparations, explicit sample-quality controls, and observables that distinguish coherent delocalization, resonant transport, or spin-sensitive effects from classical hopping or preparation artifacts Quantum in Biology Information Processing More formal or speculative at present, with the main gap being direct evidence for physiological nonclassical states or predictive advantage beyond classical alternatives Use direct cluster or state detection where possible, physiological tests of spin relaxation and coherence, and strict model-comparison benchmarks for information-processing formalisms Quantum for Biology Diamond quantum sensing Technologically mature device physics; the biological inference chain remains the limiting step Require cross-laboratory calibration, uncertainty propagation, and end-task biological validation Quantum for Biology Quantum-enhanced NMR/MRI and imaging Rapidly emerging, with several credible methodological advances and a clear benchmark structure Use dose-, time-, and sample-matched comparisons against the strongest classical modalities Quantum for Biology Quantum simulation, drug discovery, and quantum machine learning Promising but still benchmark-limited; strong claims should remain prospective Use blind targets, prospective validation, transparent compute accounting, and stronger classical baselines Biology for Quantum DNA origami and self-assembly platforms Established enabling route for positioning and fabrication Demonstrate large-scale reproducibility within integrated quantum-device workflows Biology for Quantum Biohybrid polaritons and biomolecular spin platforms Scientifically credible emerging device directions Require stability, fidelity, coherence, and reproducibility benchmarks under operating conditions Biology for Quantum Bioinspired neural and transport architectures Conceptually suggestive, but often still closer to design inspiration than to measured quantum advantage Show device- or algorithm-level gains that survive matched classical comparisons

### 27 May 2026 v2

@@ III Quantum in Biology
FIG. 2: Overview of selected topic areas in which quantum mechanical concepts are established, proposed, or debated in biological systems. The panels schematically illustrate molecular quantum chemistry, including DNA proton transfer and enzymatic hydrogen transfer; sensory transduction, including retinal photoisomerization in rhodopsin, olfactory receptor-level vibrational or electron-transfer hypotheses, and radical-pair magnetoreception; excited-state and redox processes, including photosynthetic exciton transfer in a leaf/chloroplast context, DNA charge transport, flavin photolyase repair, and mitochondrial electron transfer in a mitochondrial context; macromolecular transport, including ion-channel selectivity and microtubule photophysics; and proposed information-processing formalisms, including matrix- or operator-based representations of genetic-code structure and the Posner-cluster hypothesis. The figure is illustrative and does not imply that all topics have the same evidential status or biological relevance.
@@ III.3.4 Cellular respiration
Mitochondrial respiration transfers electrons from metabolic substrates through redox cofactors in the respiratory complexes of the inner mitochondrial membrane, coupling stepwise electron transfer to proton translocation and ATP synthesis. The most plausible quantum contributions in this context are tunneling corrections to electron- and proton-transfer steps, because these processes involve donor–acceptor distances and barriers for which quantum effects can become rate-relevant, particularly when coupled to protein reorganization, gating motions, and proton-coupled electron-transfer chemistry [165, 245, 93, 94, 95, 155, 31, 1, 158, 91, 92].
More speculative proposals have framed mitochondria as candidate quantum-biological systems, but the secure claim is narrower: local respiratory-chain chemistry can require quantum-mechanical electron and proton transfer, whereas the proton-motive force is a classical electrochemical gradient rather than a collective quantum state [175]. At larger scales, cristae ultrastructure, respiratory-supercomplex organization, fusion–fission dynamics, and inter-mitochondrial junctions shape membrane potential, diffusion, ROS production, and respiratory efficiency through mechanisms that are presently best interpreted within classical bioenergetics and cell biology [186, 187, 246]. Recent NADH-anisotropy measurements support strong mitochondrial structural directionality in living cells, but whether this alignment has quantum-relevant electrochemical or photonic consequences is not yet established [223]. Mitochondria are also plausible contributors to ultra-weak photon emission through ROS-driven excited-state chemistry, and recent reports of non-chemical signaling between isolated mitochondria are intriguing, but not yet decisive evidence for a functional quantum-optical communication channel [243, 166]. Decisive studies should therefore isolate perturbations that alter tunneling likelihood while minimally changing classical driving forces, redox landscape, and membrane potential, and then link those microscopic changes to biochemical or cellular outputs.
III.3.5 Ultra-weak photon emission
Ultra-weak photon emission (UPE), often called biophoton emission, is the spontaneous emission of very low-intensity light from biological systems. It is generally understood as an endogenous photophysical signal linked to oxidative metabolism, reactive oxygen species, and electronically excited molecular species generated during lipid, protein, and nucleic-acid oxidation [192, 167]. Recent studies support UPE as a possible non-invasive marker of biological state, including cancer-related metabolic changes and dynamic physiological responses to ischemia–reperfusion [168, 29]. Radiation-bystander studies further suggest that UPE or radiation-induced biophoton signals can participate in downstream cellular responses, including exosome-mediated effects, but these observations are best interpreted at present as photon-triggered photochemical or biochemical signaling rather than evidence for nonclassical light [138, 240].
For biological interpretation, UPE is best viewed as an integrated redox-optical signal rather than a source-specific reporter. The detected emission can combine contributions from mitochondrial respiration, lipid peroxidation, singlet-oxygen chemistry, inflammatory redox activity, tissue optical filtering, and detector response, and is therefore most informative when interpreted alongside independent metabolic, electrophysiological, and ROS measures [192, 167, 243].
For quantum biology, the key distinction is between UPE as an established photophysical observable and stronger claims about functional quantum resources. Although UPE involves quantum transitions from excited molecular states to lower-energy states, this does not by itself imply long-lived optical coherence, entanglement, nonclassical light, or biologically meaningful photon-mediated communication. Critical analyses of photocount statistics conclude that evidence for coherence or nonclassicality in biological UPE remains limited [57]. Decisive progress will require calibrated spectral, temporal, and photon-statistical measurements under controlled physiological or redox perturbations, with independent metabolic, electrophysiological, ROS, and tissue-optical validation.
III.3.6 Chiral induced spin selectivity in biological electron transfer
@@
III.3.7 Ferritin protein nanoparticles, electron tunneling, and nanomagnetism
Ferritin provides a distinct protein-nanoparticle case at the boundary between biomolecular electron transport, nanomagnetism, iron redox chemistry, and biological interpretation. Ferritin is a chiral protein cage containing an iron-oxide core, and biochemical reduction studies indicate that electron transfer to the ferritin core and iron release can occur without requiring reductant entry into the ferritin interior [250]. Ferritin-based junction and multilayer experiments have further reported long-range electron tunneling, iron-loading-dependent transport regimes, and Coulomb-blockade-like behavior under device conditions [134, 32, 202, 67]. Related work on chiral-molecule-coated superparamagnetic iron oxide nanoparticles has shown that asymmetric adsorption of chiral molecules can stabilize single-domain ferromagnetism at approximately 10 nm, with the magnetic orientation depending on molecular handedness [124]; a later study reported related magnetic-monopole-like behavior in superparamagnetic nanoparticles coated with chiral molecules [267]. These findings do not demonstrate the same effect in ferritin in vivo, but they make ferritin, because it combines a chiral protein cage with an iron-oxide core, a plausible biomolecular system in which electron transport, spin selectivity, and nanomagnetism could intersect. These results make ferritin relevant to quantum-biology discussions of electron tunneling and biomolecular nanomagnetism, but they should not yet be treated as evidence for an established in vivo signaling mechanism. Recent review-level proposals have suggested that single-particle or array-level ferritin electrical and magnetic properties could influence cellular bioelectric or biomagnetic processes [203]. At present, however, the biological claims remain hypothesis-level because device geometries, iron loading, hydration, aggregation state, redox environment, and cellular localization can all change the relevant transport and magnetic observables. Decisive tests would require in situ or near-native measurements that vary ferritin iron loading, redox state, magnetic field, and chiral or spin-sensitive readouts while linking the result to a defined biological endpoint.
@@ III.6 Section synthesis
The quantum-in-biology landscape separates less by biological domain than by evidential maturity. The most secure cases are microscopic quantum-chemical contributions that are tightly constrained by kinetics and mechanism, especially enzymatic hydrogen tunneling and related nonadiabatic proton- or electron-transfer chemistry. An intermediate band contains areas such as radical-pair magnetoreception, photosynthetic exciton and vibronic dynamics, DNA charge transport, DNA repair photochemistry, and respiration-related transfer processes, where the mechanistic frameworks are physically serious but the case for biological necessity remains conditional and perturbation-dependent. A more speculative band contains system-level extrapolations, including strong microtubule-wide physiological claims, ferritin-based bioelectric or biomagnetic signaling hypotheses, Posner-cluster neurobiology, and physical interpretations of quantum-information formalisms in genetic or neural information processing, where decisive tests are still missing or not yet sufficiently discriminating [137, 151, 102, 113, 77]. A compact topic-by-topic summary of this section is provided in Appendix Table 2.
@@ IV Quantum for Biology
FIG. 3: Overview of representative quantum-enabled and quantum-inspired tool classes for biological research. The panels schematically illustrate quantum simulation of biomolecular systems, quantum-assisted molecular design and drug discovery, quantum machine learning for bioinformatics, diamond NV sensing in biological contexts, quantum-enhanced NMR and MRI, quantum-enhanced imaging and spectroscopy, and quantum-inspired decision-theory approaches. The figure summarizes methodological directions discussed in the section and does not imply established quantum advantage for each application.
@@ V Biology for Quantum
FIG. 4: Overview of representative ways in which biological structures, materials, and design principles may inform or support quantum technologies. The panels schematically illustrate self-assembly and fabrication strategies, including DNA origami scaffolds for nanoscale placement and biotemplated conductive pathways; biohybrid excitonic platforms, including light-harvesting complexes in optical microcavities; biomolecular spin platforms, including protein-based spin sensors and qubit candidates; and bioinspired architectures, including noise-robust transport inspired by photosynthesis and quantum or quantum-inspired neural-network architectures. These examples illustrate possible routes by which biological organization or function may be translated into quantum-device concepts, with evidential strength depending on device-level benchmarks, reproducibility, and comparison with non-biological alternatives.
@@ V.1.2 Biotemplated bottom-up fabrication
Ferritin is also relevant as a protein-cage platform for quantum-adjacent device concepts, because its iron-oxide core, self-assembled shell, and reported electron-transport behavior make it a candidate biological template for nanoscale electronic, magnetic, and sensing architectures [67, 134, 32].
@@ V.4.1 Noise-robust transport and interference inspired by photosynthesis
V.4.2 Quantum biomimetics and artificial-life-inspired algorithms
Quantum biomimetics provides another bioinspired route for quantum technologies, but its status should be treated carefully. In this context, the goal is not to claim that living systems are quantum artificial-life machines, but to use selected biological behaviors, such as self-replication, mutation, interaction, inheritance, adaptation, and death, as design motifs for quantum-information protocols and controllable toy models [11, 12, 13, 136]. Proof-of-principle work has shown that such behaviors can be encoded in quantum platforms, including an implementation of a quantum artificial-life algorithm on an IBM quantum computer, while related models have explored photon-driven self-replication and links to dissipative adaptation [13, 242]. The evidential status is therefore exploratory and design-motivating: these studies extend the repertoire of bioinspired quantum algorithms and simulations, but they do not by themselves establish a biological quantum mechanism or a practical quantum advantage. Decisive progress would require benchmarked tasks in which artificial-life-inspired quantum protocols outperform strong classical artificial-life, evolutionary, or standard quantum-algorithm baselines under transparent resource accounting.
V.4.3 Brain-inspired quantum and quantum-inspired neural computation
@@ V.5 Section synthesis
Biology for quantum is strongest when the biological contribution is not merely illustrative but difficult to replace by conventional design strategies, and when that contribution is measured through explicit device, fabrication, sensing, or algorithmic benchmarks. DNA origami and related self-assembly platforms already provide concrete enabling capabilities for nanoscale positioning and fabrication. Biohybrid polaritonic systems and biomolecular spin platforms are scientifically credible emerging directions, but they still require stronger benchmarks for stability, reproducibility, and device-level performance. By contrast, quantum biomimetic and artificial-life-inspired protocols, together with broader brain-inspired or quantum-like architectures, should presently be treated as hypothesis-generating and design-motivating unless they can be tied to explicit hardware, sensing, or algorithmic advantages under matched comparisons [127, 135, 74, 2, 12, 13, 136, 3, 214]. A compact topic-by-topic summary of this section is provided in Appendix Table 4.
@@ VI Evolution and Future Perspectives
A useful discipline across this interface is to require every strong claim to answer five questions explicitly: What specific quantum resource is being invoked? What is the minimal mechanistic or algorithmic model that generates distinctive predictions? What is the strongest classical alternative under the same conditions? What perturbation, benchmark, or matched comparison would discriminate between them? And what biological or engineering output would count as success? Claims that cannot yet be mapped onto this framework are often best presented as hypotheses or design directions rather than as established results. This caution is particularly important in adjacent areas such as photobiomodulation and light-responsive mitochondrial signaling, where reported biological effects are clearly of interest, even though the full mechanistic chain remains incompletely resolved and is unlikely to be captured by a single explanatory model [208, 142, 15]. Similarly, work on ultra-weak photon emission highlights the importance of distinguishing experimentally established photophysical signals, whose intensity, spectrum, and timing may provide multidimensional redox-metabolic readouts, from stronger claims about coherence or nonclassicality, for which reliable evidence remains limited [56].
@@ VII Acknowledgments
During the preparation of this manuscript, GPT-assisted tools were used in a limited editorial role to support clarity, organization, and language refinement. Selected schematic visual elements were also generated or refined using GPT-assisted image tools, with final figures assembled and edited by the authors. All scientific content, literature selection, interpretation, figure selection, and final wording were independently reviewed and approved by the authors. The authors take full responsibility for the accuracy, integrity, and originality of the manuscript.
@@
The authors thank Chris Rourk for his thoughtful feedback and for highlighting ferritin-related literature that helped improve the discussion of biomolecular electron tunneling, nanomagnetism, and possible bioelectric or biomagnetic mechanisms.
@@ Appendix A Summary tables across the map
Direction Topic cluster Present status Most efficient upgrade path Quantum in Biology Molecular quantum chemistry and transformation Most secure where microscopic quantum-chemical effects are tightly constrained by kinetics and mechanism, especially enzymatic hydrogen tunneling and related PCET chemistry Link microscopic transfer mechanisms more tightly to full catalytic cycles, polymerase-resolved outcomes, and perturbation-resolved biological phenotypes Quantum in Biology Sensory transduction Strongest for photon detection and radical-pair spin chemistry; olfaction remains more theoretical and receptor-level evidence is limited Use receptor-, protein-, and organism-level perturbations tied directly to sensory outputs, with controls for classical binding, adaptation, and transduction effects Quantum in Biology Excited-state, redox, and spin transfer Strong physicochemical or ultrafast evidence in several systems, but functional necessity under biological conditions remains conditional and perturbation-dependent Use dephasing-, detuning-, redox-, spin-, or photochemical perturbations tied directly to yield, signaling, repair, or bioenergetic output Quantum in Biology Macromolecular transport and collective effects Experimentally and theoretically active, especially for microtubule photophysics and spin-sensitive tubulin effects, but not yet established as physiological quantum mechanisms Use independent replication, standardized preparations, explicit sample-quality controls, and observables that distinguish coherent delocalization, resonant transport, or spin-sensitive effects from classical hopping or preparation artifacts Quantum in Biology Information Processing More formal or speculative at present, with the main gap being direct evidence for physiological nonclassical states or predictive advantage beyond classical alternatives Use direct cluster or state detection where possible, physiological tests of spin relaxation and coherence, and strict model-comparison benchmarks for information-processing formalisms Quantum for Biology Diamond quantum sensing Technologically mature device physics; the biological inference chain remains the limiting step Require cross-laboratory calibration, uncertainty propagation, and end-task biological validation Quantum for Biology Quantum-enhanced NMR/MRI and imaging Rapidly emerging, with several credible methodological advances and a clear benchmark structure Use dose-, time-, and sample-matched comparisons against the strongest classical modalities Quantum for Biology Quantum simulation, drug discovery, and quantum machine learning Promising but still benchmark-limited; strong claims should remain prospective Use blind targets, prospective validation, transparent compute accounting, and stronger classical baselines Biology for Quantum DNA origami and self-assembly platforms Established enabling route for positioning and fabrication Demonstrate large-scale reproducibility within integrated quantum-device workflows Biology for Quantum Biohybrid polaritons and biomolecular spin platforms Scientifically credible emerging device directions Require stability, fidelity, coherence, and reproducibility benchmarks under operating conditions Biology for Quantum Bioinspired transport, artificial-life, and neural architectures Conceptually suggestive, but often still closer to design inspiration than to measured quantum advantage Show device- or algorithm-level gains that survive matched classical comparisons